Run Coordination Meetings: Fall 2016 Run

From GlueXWiki
Jump to: navigation, search

Meeting General Agenda

Connect Information

To connect from the outside by phone please do the following

1.) To join via a Web Browser, go to the page [1] https://bluejeans.com/660743227.

2.) To join via Polycom room system go to the IP Address: 199.48.152.152 (bjn.vc) and enter the meeting ID: 660743227.

3.) To join via phone, use one of the following numbers and the Conference ID: 660743227.

4.) The moderator code is 8394

Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada

Specific instructions for connecting from JLab's Hall D conference room:

  • Turn polycon on if necessary (do it before turning the computer on)
  • With the polycon, place a call at 199.48.152.152
  • Press # to enable the polycom keypad, then enter the meeting id: 660743227 and #
  • You may have to unmute the microphone: #*4
  • Turn the computer on if needed

Beam and Hall configuration

  • Beam energy expected: 11.632 GeV (include synchrotron radiation losses, except for the Hall D ramp).
  • Solenoid on at 1350A. Any field changes should be avoided if possible.
  • Beam current: CW 1 nA-1.5 μA. 250 MHz frequency.
  • FFB on or Position Slow Lock on (FFB doesn't operate below 50nA). Current and energy locks on?
  • Radiators:
    • Goniometer diamond radiators:
      • J1a50 (50 μm, 4.1·10-4 R.L.),
      • JD70-100 (55 μm, 4.5·10-4 R.L.),
      • JD70-105 (52 μm, 4.9·10-4 R.L.),
      • JD70-104 (23 μm, 1.6·10-4 R.L.).
    • We also have one 10 μm Al radiator.
    • Amorphous
      • 1.5 µm Al (1.7·10-5 R.L.)
      • 10 µm Al (11.2·10-5 R.L.)
      • 30 µm Al (33.7·10-5 R.L.)
  • Tagger quadrupole on (-4.2 A).
  • Collimator hole: Both 5 mm diameter and 3.4 mm diameter. Default running: 5.0 mm

General Information

This document describes the run plan for the Hall D Spring 2016 run.

  • Schedule for the Spring 2016 run, updated for the delays due to arc 7 power supply and RF separator vacuum leak. (Original dates were according to the long term calendar)
    1. Dec. 12th-13th: Electron beam restoration
    2. Dec. 14th-Dec. 21st, 8am: Hall D Fall run. Hall A will be at 4 pass with relatively low current (10 to 20 μA)
    3. Every Tuesdays: 12h of beam study. Every Thursdays, if necessary: 8h of RF recovery.
  • Goals for the Fall 2016 run were.
  1. GlueX low luminosity data taking
  2. High luminosity trigger and DAQ tests.
  3. ToF checkout.
  • Revised goals for only 1 week of running during the Fall 2016 run are.
  1. Trigger and DAQ status check.
  2. High luminosity trigger and DAQ performance study.
  3. Checkout of new diamond(s) and of goniometer (para vs perp rate difference)
  4. ToF checkout.
  5. Enough data taking for global check out.


Runplan

  1. Re-establish electron and photon beams (assume 3 days).
    • Ion Chambers trip threshold must be set.
    • Tune electron beam with radiator retracted
      • Amorphous Radiator (and diamond) must be retracted. Collimator should be in blocking position. CDC/FDC are off
      • Radiation levels should be closely monitored by MCC crew.
        • Typical levels are (for radiator retracted, collimator fully blocking, electron beam current 50 nA)
          • RAD102_P1 ~ 3 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2 ~ 15 mrad/hr (tagger area, gammas, near electronics racks).
          • RAD102_P3 ~ 0.1 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1 < 0.1 mrad/hr (collimator cave, gammas).
          • RAD508_P2 < 0. mrad/hr (Hall. Neutrons).
      • Hall D tagger CARMS, Ion Chamber and BLM threshold for beam trips should not be changed without approval of Hall D leader.
    • Check bleedthrough level. 10min Hovanes. Hall A needs to be running
      • Beam magnetic separation during Oct. We expect no significant bleedthrough.
    • Establish good photon beam
    • Make sure tagger quadrupole is on (-4.2 A).
      • Insert 2*10-5 RL radiator. (Call MCC and ask to mask the FSD each time the radiator is moved)
        • Typical levels are (for 2*10-5 radiator, electron beam current 50 nA)
          • RAD102_P1= 2-10 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2= 40-50 mrad/hr (tagger area, gammas, near electronics racks)
          • RAD102_P3= ~2 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1= ~10 mrad/hr (collimator cave, gammas).
          • RAD508_P2 < ~ 0. mrad/hr (Hall. Neutrons).
    • Once radiation levels are good, proceed with harp scans (Hovanes/Todd. 15 min)
      • Take tagger harp scans. Check consistency with Accelerator's harps. Study beam focusing
    • Radiation level study (1h):
      • Run 5 min without radiator (Call MCC and ask to mask the FSD each time the radiator is moved). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 5 min with the 10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 3 min with the 3*10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
    • Active Collimator calibration. (20 + 20 mins)
      • re-Insert the 10-4 radiator (Call MCC). Use about 200 nA beam current. (This corresponds to the expected photon flux for standard production runs).
      • Determine maximum photon beam transmission by doing 2D scans (x: beam scan, y: beam scan. ±5mm with 0.5mm steps. Stay 5 seconds with stable beam for each steps Remind MCC that they should use only the last corrector (5C11A) to adjust the beam position. If the adjustment results in unacceptable e- beam position in the tagger dump, remind MCC that they should use the tagger magnet only to readjust the e- beam position in the tagger dump. Meanwhile, monitor the rate at the active target and rates (including coinc.) in the pair spectrometer.
      • Repeat procedure with 2*10-5 radiator (Call MCC) and 50 nA beam current.
      • In addition do x/y scans using the collimator x/y motions and check consistency (Hovanes. 1h)
      • Run fast calibration procedure with beam rastering
      • Record on the white board what is the profiler position corresponding to the center of the A.C.
    • Scan the photon beam profile with PS rates and collimator x/y motions Need to have procedure/code ready. Do we need to install the collimator plug?? (Hovanes ??h)
    • Verify with MCC that position slow lock is back on (locking on profiler and for values corresponding to optimal photon flux).
  1. Detector and beamline checkout.
    • L1 trigger tuning. Check new firmware (Alex. Somov. 3h). These tests will be continued later on and completed (3h+4h).
    • Reestablish DAQ (Sergei, 3h). Sergey to establish optimal beam current for DAQ (with 10-4 radiator). These tests will be completed later (3h).
    • General detector calibration (Sean Dobbs . 4h). Use 10-4 radiator. Call Sean when getting ready for this task (at any hour): (703) 887-1776.
      • At the same time, do Hodoscope voltage scans (Nathan S.)
      • Run 15 min in Long Mode.
      • Push DAQ performances by running with 15 min 15% above optimal beam current (as determined by Sergey above). Then run another 15 min 50% above.
  2. Realign old 50 μm diamond (J1a50). Later on, align at least one new 50 μm diamond (Hovanes, Paul Mattione. 12h/diamond).
    • ToF tests (Benedikt Zihlmann. 6h).
      • At the same time, do CDC voltage scans.
      • At the same time, do PS magnetic field scans. (Alex. Somov. 2h)
  3. L1 trigger test for high-rate performance tests (Alex Somov. 8h+8h. Use up to 1 μA. DC off if necessary)
    • Check rates for various beam currents and thr. settings.
  4. Run PS at lower B-field so that the TPol can check the low photon energy tail polarization. PS set-up so that it sees 3 GeV photons (Alex Somov/Nathan Sparks/Mike Dugger 4h/maybe parasitic?)
  5. DAQ high-rate performance tests (Sergei Furletov 2h)
  6. high-current mapping (Sergei Furletov/Alex Somov/David Lawrence 2h) plan
  7. L3 high-rate performance tests (David Lawrence 4h, 2h+2h, separated by a few days) plan
  8. Check bleedthrough level once Hall A is at 5th pass and RF separator is used (Nov. 1st) Hovanes. Hall A needs to be running
    • Need to assess level of acceptable bleedthrough
  9. Gluex data production
    • Tagger quadrupole on
    • Harp scans once a day: 15 min.
    • Empty target run every week (2h: 1- h to empty/fill back the target, 1h running). Standard production current.
    • Amorphous run. Every day. 1h Standard production current. Time to switch from pol. to unpol.: 10min. time to switch from unpol. to pol.: 30min.
    • TAC runs. 2 TAC runs (one at the beginning, one in the middle of the run). (Hovanes, Alex Somov, 5h: 4h run with thick converter. 1- h target empty+fill)
    • Physics production data with diamonds and 5 mm hole. Ratio of para./transverse data:50/50. 2h runs and no more, as it may be hard to correct for calibration drifts. Switch polarization every run. Time to switch: 10min. DAQ start run overhead: 3min.
      • Start run with old 50 μm diamond (J1a50). When new 20 μm diamonds is aligned, assess its quality and decide what diamond should be used. Possibly some running on new 50 μm diamond is its quality warrants it.
      • Luminosity:
        • Start at about 100 nA.
        • Then, once flux calibration is available from TAC, run at 107 photon flux.
        • Then, assess if we can run at a higher flux. (Paul M., to establish criteria of what flux we will be comfortable with, using e.g. accidental rate and/or detector aging).
      • Trigger configuration: First week: run in the same conditions as Spring 16. Then, evaluate if we can run at higher thr.
  10. Parasitic run for PrimEx detector: move ComCal close to the beam line and check rates.
  11. Parasitic TRD run

Hall D white board (current run plan)

Hall D counting room white board

Runplan time charts



Accelerator status

Hot Checkout

  1. Hot Check OUT Status 100% ready ; 0% checked ; 0% not ready

Readiness for Fall 16 run

  • <Solenoid (Eugene/Beni) is Ready.
    • Check with Spring16 data that running at 1350A vs 1200A is beneficial (Paul Mattione/Mike/Alex) Running at 1350A is prefereable for GlueX.
    • Will ramp to 100 A a few days before the start of the run, then leave power supply on at 0 V until experiment. Done.
  • Surveys (Tim/Tom)
    • Target Fiducialization/alignment (Monday Sept. 19) Done
    • Tagger Microscope alignment, Tagger Hodoscope Survey (Tuesday Sept 20) Done
    • Start Counter Alignment, Target Cart Alignment (Thu. Sept 22) Done
    • Primary Collimator Fiducialization/alignment, Beam Profiler Alignment (Fri Sept 23) Done
    • CompCal Fiducialization/Alignment (Monday Sept 26) Done
    • PS High Granularity Alignment, PS beamline alignment (Tuesday Sept 27) Done
    • TAGH needs to be resurveyed.
    • Goniometer fiducialization.
  • DAQ (Sergei)
    • Waiting on Trigger to be ready.
    • Most systems are tested. Successfully ran at 60 kHz, 99% livetime.
  • Trigger (Alex S.)
    • System is robust. Running on cosmics. Still a few problems to be addressed:
      • FDC F125 send error messages. Fixed
      • F250 DAC sometimes fails. Fixed
      • Problem with PS/Start Counter crate: TDC sometimes generates readout problem. Problem seems to now be gone but need to be checked.
      • Problem with TagM. TDC crate. Fixed now.
  • Beam line
    • Diamond status.
      • 3 new and characterized diamonds are now JLab.
    • Tagger hodoscope (Nathan/Franz)
      • Remove protective plastic tubes. Franz to remove part of them this week (Sept. 19-23). Done on 19th.
      • Once tubes are removed, need to check for light leak with scalers. Done.
      • Needs to be resurveyed
      • Needs testing with DAQ<strike>
    • <strike>Tagger microscope (Alex B./Richard) TAGM is ready
      • waiting for survey to be completed before doing hot check out. Done
      • Need to put back shielding.
      • 1 MPOD to power preamp errors out, Alex B following up. Blower on and shroud covered, ready for dark rate testing.
    • Goniometer (Hovanes)
      • Goniometer is back. Issue with control hardware. Need to be fixed and tested.
      • To be fiducialized. Done.
    • Collimator, Active collimator, profiler, flux monitor (Alex B./Hovanes).
      • AC installed and wired, needs DAQ testing A.C. needs some fine tuning.
      • Flux monitor cabled (was this verified?)
      • Collimator vertical motion control is done. 2D scan procedure is written.
      • Profiler is ready.
    • PS system (Alex S.)
      • Need software for online monitoring of photon flux.
      • Replace thermal pads?: Will decide this week (Sept. 19-23). No time to installed them for this run.
      • Installed, alignment in progress
      • High Granularity detector had Power Supply glitches. Now fixed.
      • Needs DAQ testing
    • Polarimeter (M. Dugger/Nathan) Polarimeter is ready
      • Now under vacuum.
      • Signals seen on all channels with source and scope.
      • DAQ tested
    • Target (Tim/Alexandre D.)
      • Ice test done. Leak fixed. Thin frost seen. Will add MLI to improve insulation. Leak fixed.
      • Target is now inserted.
    • TAC (Hovanes) TAC is ready and (re)checked.
    • Slow control (Hovanes)
      • Goniometer motion.
      • Need to work on vertical col. motion and need to program ComCal. (PrimEx detector) motion.
      • Have implemented temperature thr. that shut down the crates. For now, they will operate between 62F<T<82F
    • RF (Paul Mattione)
      • Fernando is implementing new TDC firmware for CAEN.
      • New board is installed.
      • Then, Paul to test the RF system.
    • Radiation monitors
      • Need to move a photon probe near goniometer. Done
      • Install OSL and chromosticks<strike>
  • Detector systems
    • ST (Mark Ito)
      • <strike>Was test-cabled. Will be de-cabled, inserted and re-cabled.
      • Inserted, powered up, working with stand alone DAQ.
      • Crate burned 2 TI modules. Removed from main DAQ readout.
      • ST is now re-installed and need to be rechecked..
    • FDC (Lubomir)
      • One sector doesn't hold HV (no change)
      • Strange noise seen on each 3rd card of each chamber. Seen at 60 kHz but not at 1 kHz.
    • CDC (Beni)
      • Up and running
      • Calibrated baselines, problems with 2 channels
      • 2 weekends of cosmic data desired.
    • BCal (Elton)
      • Was turned on on Sept. 14th. Turned on LED pulser too. One of the LED low voltage doesn't work. Chris is checking it.
      • Stand alone DAQ run, all LV and SiPMs working. Intermittent problems with ADCs, possibly initialization.
      • Checked data and saw that many channels are missing. TDC were missing. This is fixed now
    • FCal (Adesh)
      • Was turned on this week (Sept. 19-23).. On since then.
      • Replaced 22 bases week of 19-24 Sept, more failures since at rate of 1-2 per day.
      • Keep replacing basis when needed.
      • Adesh to do the HCO. Done
    • TOF (Mark Ito/Beni/Alex. Ostrovidov)
      • Beni: was able to do a first calibration but unable to do it anymore. Replacing a HV module.
      • Currently taking data for TDC correction tables.
      • Paul E. would like to install a lucite shield. Fitting it was tested and it fits fine.
      • Need to verify that all ADC channels are set-up correctly.
  • Online monitoring (David L.)
    • Implemented Root 6
    • Check compatibility with new F250 format. (One aspect of format might cahnge this week)
    • Ready to go with occupancy and new higher level reconstruction histograms
    • As ready as can be.
  • Calibration/reconstruction/offline monitoring (Paul Mattione)
    • Ready.
  • Level 3 Trigger (David L.)
    • As ready as can be given the uncertainty on what to do.
  • Safety documents (Beni):
    • ESAD must be updated to include safety procedure for ComCal. motion.
    • ESAD ready, RSAD should be ready.
  • Hall Status/safety (Tom/Tim)
    • Ready
  • Other topics?